The Synthesis Step: From Ill-posed Problems to PFD

For every complicated problem, there is an answer that is short, simple and wrong

H.L. Mencken Given: An ill-posed "word" problem To develop a PFD which includes the following information:

Unit operations Mass balances (flow rates in and out of each unit operation) Energy balances (heat requirements, enthalpies) Operating temperature and pressure of each unit operation Equipment size Equipment cost

This information is necessary for Feasibility analysis Detailed profitability analysis Generating P&IDs for construction

The Synthesis Step: From Ill-posed Problems to PFD – p.3/32

Given: An ill-posed "word" problem To develop a PFD which includes the following information:

Unit operations Mass balances (flow rates in and out of each unit operation) Energy balances (heat requirements, enthalpies) Operating temperature and pressure of each unit operation Equipment size Equipment cost

This information is necessary for Feasibility analysis Detailed profitability analysis Generating P&IDs for construction

The Synthesis Step: From Ill-posed Problems to PFD – p.3/32

Given: An ill-posed "word" problem To develop a PFD which includes the following information:

Unit operations Mass balances (flow rates in and out of each unit operation) Energy balances (heat requirements, enthalpies) Operating temperature and pressure of each unit operation Equipment size Equipment cost

This information is necessary for Feasibility analysis Detailed profitability analysis Generating P&IDs for construction

The Synthesis Step: From Ill-posed Problems to PFD – p.3/32

Given: An ill-posed "word" problem To develop a PFD which includes the following information:

Unit operations Mass balances (flow rates in and out of each unit operation) Energy balances (heat requirements, enthalpies) Operating temperature and pressure of each unit operation Equipment size Equipment cost

This information is necessary for Feasibility analysis Detailed profitability analysis Generating P&IDs for construction

The Synthesis Step: From Ill-posed Problems to PFD – p.3/32

Given: An ill-posed "word" problem To develop a PFD which includes the following information:

Unit operations Mass balances (flow rates in and out of each unit operation) Energy balances (heat requirements, enthalpies) Operating temperature and pressure of each unit operation Equipment size Equipment cost

This information is necessary for Feasibility analysis Detailed profitability analysis Generating P&IDs for construction

The Synthesis Step: From Ill-posed Problems to PFD – p.3/32

Given: An ill-posed "word" problem To develop a PFD which includes the following information:

Unit operations Mass balances (flow rates in and out of each unit operation) Energy balances (heat requirements, enthalpies) Operating temperature and pressure of each unit operation Equipment size Equipment cost

This information is necessary for Feasibility analysis Detailed profitability analysis Generating P&IDs for construction

The Synthesis Step: From Ill-posed Problems to PFD – p.3/32

Given: An ill-posed "word" problem To develop a PFD which includes the following information:

Unit operations Mass balances (flow rates in and out of each unit operation) Energy balances (heat requirements, enthalpies) Operating temperature and pressure of each unit operation Equipment size Equipment cost

This information is necessary for Feasibility analysis Detailed profitability analysis Generating P&IDs for construction

The Synthesis Step: From Ill-posed Problems to PFD – p.3/32

Basic Steps in Flowsheet Synthesis

Gather Information

Gather information about the process chemistry Generate flow diagram based on Douglas Hierarchy Solve mass and energy balances Estimate equipment size based on flow rates from previous step Estimate equipment cost based on size from previous step Optimize process Gather information about the process chemistry Generate flow diagram based on Douglas Hierarchy Solve mass and energy balances Estimate equipment size based on flow rates from previous step Estimate equipment cost based on size from previous step Optimize process Gather information about the process chemistry Generate flow diagram based on Douglas Hierarchy Solve mass and energy balances Estimate equipment size based on flow rates from previous step Estimate equipment cost based on size from previous step Optimize process Gather information about the process chemistry Generate flow diagram based on Douglas Hierarchy Solve mass and energy balances Estimate equipment size based on flow rates from previous step Estimate equipment cost based on size from previous step Optimize process Gather information about the process chemistry Generate flow diagram based on Douglas Hierarchy Solve mass and energy balances Estimate equipment size based on flow rates from previous step Estimate equipment cost based on size from previous step Optimize process Gather information about the process chemistry Generate flow diagram based on Douglas Hierarchy Solve mass and energy balances Estimate equipment size based on flow rates from previous step Estimate equipment cost based on size from previous step Optimize process Patent Literature

US Patent Office

Handbooks and Encyclopediae Encyclopedia of Chemical Technology, Kirk and Othmer Perry’s Chemical Engineers’ Handbook

Online Data Bases NIST Database DECHMA Database Patent Literature

US Patent Office

Handbooks and Encyclopediae Encyclopedia of Chemical Technology, Kirk and Othmer Perry’s Chemical Engineers’ Handbook

Online Data Bases NIST Database DECHMA Database Patent Literature

US Patent Office Handbooks and Encyclopediae

Encyclopedia of Chemical Technology, Kirk and Othmer Perry’s Chemical Engineers’ Handbook

Online Data Bases NIST Database DECHMA Database Patent Literature

US Patent Office Handbooks and Encyclopediae

Encyclopedia of Chemical Technology, Kirk and Othmer Perry’s Chemical Engineers’ Handbook

Online Data Bases NIST Database DECHMA Database

Case Study: Ethanol Synthesis

\underline{Main Reaction}:

CH_{2} = CH_{2} + H_{2}O → CH_{3}CH_{2}OH

5-7 % conversion

\underline{Other reactions}:

2CH_{3}CH_{2}OH ⇋ C_{2}H_{5} − O − C_{2}H_{5} + H_{2}O

Equilibrium reaction

C_{3}H_{6} + H_{2}O → CH_{3}CH_{2}CH_{2}OH

0.5-0.7 % conversion

Trace amounts of croton aldehyde are formed.

The Synthesis Step: From Ill-posed Problems to PFD – p.6/32

Case Study: Ethanol Synthesis

\underline{Main Reaction}:

CH_{2} = CH_{2} + H_{2}O → CH_{3}CH_{2}OH

5-7 % conversion

\underline{Other reactions}:

2CH_{3}CH_{2}OH ⇋ C_{2}H_{5} − O − C_{2}H_{5} + H_{2}O

Equilibrium reaction

C_{3}H_{6} + H_{2}O → CH_{3}CH_{2}CH_{2}OH

0.5-0.7 % conversion

Trace amounts of croton aldehyde are formed.

The Synthesis Step: From Ill-posed Problems to PFD – p.6/32

Case Study: Ethanol Synthesis

\underline{Main Reaction}:

CH_{2} = CH_{2} + H_{2}O → CH_{3}CH_{2}OH

5-7 % conversion

\underline{Other reactions}:

2CH_{3}CH_{2}OH ⇋ C_{2}H_{5} − O − C_{2}H_{5} + H_{2}O

Equilibrium reaction

C_{3}H_{6} + H_{2}O → CH_{3}CH_{2}CH_{2}OH

0.5-0.7 % conversion

Trace amounts of croton aldehyde are formed.

The Synthesis Step: From Ill-posed Problems to PFD – p.6/32

Case Study: Ethanol Synthesis

\underline{Main Reaction}:

CH_{2} = CH_{2} + H_{2}O → CH_{3}CH_{2}OH

5-7 % conversion

\underline{Other reactions}:

2CH_{3}CH_{2}OH ⇋ C_{2}H_{5} − O − C_{2}H_{5} + H_{2}O

Equilibrium reaction

C_{3}H_{6} + H_{2}O → CH_{3}CH_{2}CH_{2}OH

0.5-0.7 % conversion

Trace amounts of croton aldehyde are formed.

The Synthesis Step: From Ill-posed Problems to PFD – p.6/32

Case Study: Ethanol Synthesis

\underline{Main Reaction}:

CH_{2} = CH_{2} + H_{2}O → CH_{3}CH_{2}OH

5-7 % conversion

\underline{Other reactions}:

2CH_{3}CH_{2}OH ⇋ C_{2}H_{5} − O − C_{2}H_{5} + H_{2}O

Equilibrium reaction

C_{3}H_{6} + H_{2}O → CH_{3}CH_{2}CH_{2}OH

0.5-0.7 % conversion

Trace amounts of croton aldehyde are formed.

The Synthesis Step: From Ill-posed Problems to PFD – p.6/32

Available process High temperature (535K − 575K) High pressure (68 atm) Homogeneous, noncatalytic reactor

\underline{Coking considerations} : Chemistry department advises that mole fraction of methane in the reactor feed should be less than 10% to prevent coking.

\underline{Equilibrium considerations}: Excess water will push the equilibrium towards ethanol.

Available process High temperature (535K − 575K) High pressure (68 atm) Homogeneous, noncatalytic reactor

\underline{Coking considerations} : Chemistry department advises that mole fraction of methane in the reactor feed should be less than 10% to prevent coking.

\underline{Equilibrium considerations}: Excess water will push the equilibrium towards ethanol.

Available process High temperature (535K − 575K) High pressure (68 atm) Homogeneous, noncatalytic reactor

\underline{Coking considerations} : Chemistry department advises that mole fraction of methane in the reactor feed should be less than 10% to prevent coking.

\underline{Equilibrium considerations}: Excess water will push the equilibrium towards ethanol.

\underline{Physical Properties}:

At 1 atm, methane, ethylene and propylene boil at very low temperatures.

It is not possible to condense methane and ethylene at room temperature.

Propylene has a vapor pressure of 15 atm at 310K. => expensive distillation At 1 atm, methane, ethylene and propylene boil at very low temperatures.

It is not possible to condense methane and ethylene at room temperature.

Propylene has a vapor pressure of 15 atm at 310K. => expensive distillation At 1 atm, methane, ethylene and propylene boil at very low temperatures.

It is not possible to condense methane and ethylene at room temperature.

Generate Flow Diagram

Propylene has a vapor pressure of 15 atm at 310K. => expensive distillation ✓Gather information about the process chemistry Generate flow diagram based on Douglas Hierarchy Solve mass and energy balances Estimate equipment size based on flow rates from previous step Estimate equipment cost based on size from previous step Optimize process The Douglas Hierarchy is a formal procedure for developing flow sheets and consists of the following steps:

Step 1: Batch vs. Continuous Process Step 2: Input-Output Structure of the Process Step 3: Recycle Structure of the Process Step 4: General Structure of Separation System Step 5: Heat Exchanger Network or Process Energy Recovery System

This procedure works for a large number of chemical, biochemical, environmental, and materials processes.

The Douglas Hierarchy is a formal procedure for developing flow sheets and consists of the following steps:

Step 1: Batch vs. Continuous Process Step 2: Input-Output Structure of the Process Step 3: Recycle Structure of the Process Step 4: General Structure of Separation System Step 5: Heat Exchanger Network or Process Energy Recovery System

This procedure works for a large number of chemical, biochemical, environmental, and materials processes.

The Douglas Hierarchy is a formal procedure for developing flow sheets and consists of the following steps:

Step 1: Batch vs. Continuous Process Step 2: Input-Output Structure of the Process Step 3: Recycle Structure of the Process Step 4: General Structure of Separation System Step 5: Heat Exchanger Network or Process Energy Recovery System

This procedure works for a large number of chemical, biochemical, environmental, and materials processes.

The Douglas Hierarchy is a formal procedure for developing flow sheets and consists of the following steps:

Step 1: Batch vs. Continuous Process Step 2: Input-Output Structure of the Process Step 3: Recycle Structure of the Process Step 4: General Structure of Separation System Step 5: Heat Exchanger Network or Process Energy Recovery System

This procedure works for a large number of chemical, biochemical, environmental, and materials processes.

The Douglas Hierarchy is a formal procedure for developing flow sheets and consists of the following steps:

Step 1: Batch vs. Continuous Process Step 2: Input-Output Structure of the Process Step 3: Recycle Structure of the Process Step 4: General Structure of Separation System Step 5: Heat Exchanger Network or Process Energy Recovery System

This procedure works for a large number of chemical, biochemical, environmental, and materials processes.

The Douglas Hierarchy is a formal procedure for developing flow sheets and consists of the following steps:

Step 1: Batch vs. Continuous Process Step 2: Input-Output Structure of the Process Step 3: Recycle Structure of the Process Step 4: General Structure of Separation System Step 5: Heat Exchanger Network or Process Energy Recovery System

This procedure works for a large number of chemical, biochemical, environmental, and materials processes.

The Douglas Hierarchy is a formal procedure for developing flow sheets and consists of the following steps:

Step 1: Batch vs. Continuous Process Step 2: Input-Output Structure of the Process Step 3: Recycle Structure of the Process Step 4: General Structure of Separation System Step 5: Heat Exchanger Network or Process Energy Recovery System

This procedure works for a large number of chemical, biochemical, environmental, and materials processes.

The Douglas Hierarchy is a formal procedure for developing flow sheets and consists of the following steps:

Step 1: Batch vs. Continuous Process Step 2: Input-Output Structure of the Process Step 3: Recycle Structure of the Process Step 4: General Structure of Separation System Step 5: Heat Exchanger Network or Process Energy Recovery System

This procedure works for a large number of chemical, biochemical, environmental, and materials processes.

Step 1: Batch vs. Continuous Process

When is batch process preferred? Small quantity ( < 500 tonnes/yr), high quality, highly regulated New product, safety, sterility considerations Hard to handle chemicals and processes Large variation in feed, product and operating conditions

When is continuous process preferred? Large quantity ( > 500 tonnes/yr) Well defined process Small variation in feed, product and operating conditions

The Synthesis Step: From Ill-posed Problems to PFD – p.11/32

When is batch process preferred? Small quantity ( < 500 tonnes/yr), high quality, highly regulated New product, safety, sterility considerations Hard to handle chemicals and processes Large variation in feed, product and operating conditions

When is continuous process preferred? Large quantity ( > 500 tonnes/yr) Well defined process Small variation in feed, product and operating conditions

The Synthesis Step: From Ill-posed Problems to PFD – p.11/32

When is batch process preferred? Small quantity ( < 500 tonnes/yr), high quality, highly regulated New product, safety, sterility considerations Hard to handle chemicals and processes Large variation in feed, product and operating conditions

When is continuous process preferred? Large quantity ( > 500 tonnes/yr) Well defined process Small variation in feed, product and operating conditions

The Synthesis Step: From Ill-posed Problems to PFD – p.11/32

When is batch process preferred? Small quantity ( < 500 tonnes/yr), high quality, highly regulated New product, safety, sterility considerations Hard to handle chemicals and processes Large variation in feed, product and operating conditions

When is continuous process preferred? Large quantity ( > 500 tonnes/yr) Well defined process Small variation in feed, product and operating conditions

The Synthesis Step: From Ill-posed Problems to PFD – p.11/32

Step 2: Input-Output Structure

Process Concept Diagram

Basic economic analysis on profit margin. What chemical components enter as feed and leave as product. All reactions, desired and undesired.

The Synthesis Step: From Ill-posed Problems to PFD – p.12/32

Basic economic analysis on profit margin. What chemical components enter as feed and leave as product. All reactions, desired and undesired.

The Synthesis Step: From Ill-posed Problems to PFD – p.12/32

Basic economic analysis on profit margin. What chemical components enter as feed and leave as product. All reactions, desired and undesired.

The Synthesis Step: From Ill-posed Problems to PFD – p.12/32

Generic Block Flow Diagram A vast majority of processes can be represented by this generic BFD Feed Preparation System: This system is required to adjust T, P, and composition in preparation for the reactor.

If impurities are less than 10-20%, and these impurities do not react to form by-products, do not purify the feed.

If separation of feed impurities is difficult (e.g. azeotropes are formed), do not purify the feed.

If impurities foul the reactor catalyst or form hard-to-separate unwanted products, purify the feed.

Add inert material to feed to control exothermic rxn.

Add inert material to feed for favorable equilibrium rxn.

Feed Preparation System: This system is required to adjust T, P, and composition in preparation for the reactor.

If impurities are less than 10-20%, and these impurities do not react to form by-products, do not purify the feed.

If separation of feed impurities is difficult (e.g. azeotropes are formed), do not purify the feed.

If impurities foul the reactor catalyst or form hard-to-separate unwanted products, purify the feed.

Add inert material to feed to control exothermic rxn.

Add inert material to feed for favorable equilibrium rxn.

Case Study: Ethanol Synthesis

Alternative 1 Which alternative would you choose?

Reaction System: In this system, composition changes due to chemical reaction.

Temperature and pressure ranges are set by reaction chemistry.

Catalyst used is set by reaction chemistry.

Reaction type (e.g. packed bed, fluidized bed, CSTR) are set by reactions occuring and reactor conditions (e.g. exothermic, endothermic)

Reaction System: In this system, composition changes due to chemical reaction.

Temperature and pressure ranges are set by reaction chemistry.

Catalyst used is set by reaction chemistry.

Reaction type (e.g. packed bed, fluidized bed, CSTR) are set by reactions occuring and reactor conditions (e.g. exothermic, endothermic)

Separation System: In this system, the desired products are separated from by-products and unwanted products. The product stream may be further processed due to environmental, safety, or economic considerations.

This system is analyzed in detail in Step 4 Separation System: In this system, the desired products are separated from by-products and unwanted products. The product stream may be further processed due to environmental, safety, or economic considerations.

Step 3: Input-Output Structure

This system is analyzed in detail in Step 4 Unused reactants are recycled when:

Conversion is low. Raw material costs form a significant portion of operating costs. Unused reactants are easy to separate. Reactants are environmentally unfriendly.

Unused reactants are recycled when: Conversion is low. Raw material costs form a significant portion of operating costs. Unused reactants are easy to separate. Reactants are environmentally unfriendly.

Unused reactants are recycled when: Conversion is low. Raw material costs form a significant portion of operating costs. Unused reactants are easy to separate. Reactants are environmentally unfriendly.

Unused reactants are recycled when: Conversion is low. Raw material costs form a significant portion of operating costs. Unused reactants are easy to separate. Reactants are environmentally unfriendly.

Three ways to recycle unused reactants Separate and purify unreacted feed material from products and then recycle the reactants. Recycle feed and product together and use purge stream. Recycle feed and product together and do not use purge stream.

Three ways to recycle unused reactants Separate and purify unreacted feed material from products and then recycle the reactants. Recycle feed and product together and use purge stream. Recycle feed and product together and do not use purge stream.

Three ways to recycle unused reactants Separate and purify unreacted feed material from products and then recycle the reactants. Recycle feed and product together and use purge stream. Recycle feed and product together and do not use purge stream.

Recycle in Ethanol Synthesis

Alternative 1

How to separate EL and W from the reactor outlet?

How to separate EL and W from the reactor outlet?

How to separate EL and W from the reactor outlet?

Recycle in Ethanol Synthesis

Alternative 2

What should be the purge fraction? What happens to the DEE reaction which is an equilibrium reaction?

Alternative 2

What should be the purge fraction? What happens to the DEE reaction which is an equilibrium reaction?

Alternative 2

What should be the purge fraction? What happens to the DEE reaction which is an equilibrium reaction?

At \underline{steady state}, DEE is at equilibrium and no reactant is consumed.

The Synthesis Step: From Ill-posed Problems to PFD – p.24/32

At \underline{steady state}, DEE is at equilibrium and no reactant is consumed.

The Synthesis Step: From Ill-posed Problems to PFD – p.24/32

At \underline{steady state}, DEE is at equilibrium and no reactant is consumed.

The Synthesis Step: From Ill-posed Problems to PFD – p.24/32

Step 4: Separation System Synthesis

Flash Distillation Gas Absorption Extraction Filtration Chromatography Centrifugation Membrane Separation Crystalization Drying . .

Flash Distillation Gas Absorption Extraction Filtration Chromatography Centrifugation Membrane Separation Crystalization Drying . .

Flash Distillation Gas Absorption Extraction Filtration Chromatography Centrifugation Membrane Separation Crystalization Drying . .

Flash Distillation Gas Absorption Extraction Filtration Chromatography Centrifugation Membrane Separation Crystalization Drying . .

Guidelines for Choosing Separation Units

Use distillation when purity of both products is required.

Use gas absorption to remove one trace component from a gas stream.

Use membrane seperation to separate gases of cryogenic boiling points and relatively small flowrates.

Choose an alternative to distillation if the boiling points are very close or if the heats of vaporization are very high.

Use crystallization to purify a solid from a liquid solution.

Use extraction to purify a liquid from another liquid.

Please see Table 10.1 (page 360, TBWS) The Synthesis Step: From Ill-posed Problems to PFD – p.26/32

Use distillation when purity of both products is required.

Use gas absorption to remove one trace component from a gas stream.

Use membrane seperation to separate gases of cryogenic boiling points and relatively small flowrates.

Choose an alternative to distillation if the boiling points are very close or if the heats of vaporization are very high.

Use crystallization to purify a solid from a liquid solution.

Use extraction to purify a liquid from another liquid.

Use distillation when purity of both products is required.

Use gas absorption to remove one trace component from a gas stream.

Use membrane seperation to separate gases of cryogenic boiling points and relatively small flowrates.

Choose an alternative to distillation if the boiling points are very close or if the heats of vaporization are very high.

Use crystallization to purify a solid from a liquid solution.

Use extraction to purify a liquid from another liquid.

Use distillation when purity of both products is required.

Use gas absorption to remove one trace component from a gas stream.

Use membrane seperation to separate gases of cryogenic boiling points and relatively small flowrates.

Choose an alternative to distillation if the boiling points are very close or if the heats of vaporization are very high.

Use crystallization to purify a solid from a liquid solution.

Use extraction to purify a liquid from another liquid.

Use distillation when purity of both products is required.

Use gas absorption to remove one trace component from a gas stream.

Use membrane seperation to separate gases of cryogenic boiling points and relatively small flowrates.

Choose an alternative to distillation if the boiling points are very close or if the heats of vaporization are very high.

Use crystallization to purify a solid from a liquid solution.

Use extraction to purify a liquid from another liquid.

Use distillation when purity of both products is required.

Use gas absorption to remove one trace component from a gas stream.

Use membrane seperation to separate gases of cryogenic boiling points and relatively small flowrates.

Choose an alternative to distillation if the boiling points are very close or if the heats of vaporization are very high.

Use crystallization to purify a solid from a liquid solution.

Use extraction to purify a liquid from another liquid.

Please see Table 10.1 (page 360, TBWS)

Use distillation when purity of both products is required.

Use gas absorption to remove one trace component from a gas stream.

Use membrane seperation to separate gases of cryogenic boiling points and relatively small flowrates.

Choose an alternative to distillation if the boiling points are very close or if the heats of vaporization are very high.

Use crystallization to purify a solid from a liquid solution.

Use extraction to purify a liquid from another liquid.

Please see Table 10.1 (page 360, TBWS)

Sequencing Separation Units

Remove the largest product stream first. For distillation, remove the product with the highest heat of vaporization first. Do not recombine separated streams. Do the easy separation first. Do not overpurify streams based on their use. Remove hazardous or corrosive materials first.

Remove the largest product stream first. For distillation, remove the product with the highest heat of vaporization first. Do not recombine separated streams. Do the easy separation first. Do not overpurify streams based on their use. Remove hazardous or corrosive materials first.

Remove the largest product stream first. For distillation, remove the product with the highest heat of vaporization first. Do not recombine separated streams. Do the easy separation first. Do not overpurify streams based on their use. Remove hazardous or corrosive materials first.

Remove the largest product stream first. For distillation, remove the product with the highest heat of vaporization first. Do not recombine separated streams. Do the easy separation first. Do not overpurify streams based on their use. Remove hazardous or corrosive materials first.

Remove the largest product stream first. For distillation, remove the product with the highest heat of vaporization first. Do not recombine separated streams. Do the easy separation first. Do not overpurify streams based on their use. Remove hazardous or corrosive materials first.

Remove the largest product stream first. For distillation, remove the product with the highest heat of vaporization first. Do not recombine separated streams. Do the easy separation first. Do not overpurify streams based on their use. Remove hazardous or corrosive materials first.

Separation System for Ethanol Synthesis

Separate the reactor products into a liquid phase and vapor phase.

Design vapor recovery system and liquid recovery system separately.

Separate the reactor products into a liquid phase and vapor phase.

Design vapor recovery system and liquid recovery system separately.

Separate the reactor products into a liquid phase and vapor phase.

Design vapor recovery system and liquid recovery system separately.

Flash the reactor products to get a vapor stream and liquid stream.

The vapor stream is mainly EL, PL, and M and is recycled back.

Some EL will leave in the vapor stream which can be recovered via an absorber.

The Synthesis Step: From Ill-posed Problems to PFD – p.29/32

Flash the reactor products to get a vapor stream and liquid stream.

The vapor stream is mainly EL, PL, and M and is recycled back.

Some EL will leave in the vapor stream which can be recovered via an absorber.

The Synthesis Step: From Ill-posed Problems to PFD – p.29/32

Flash the reactor products to get a vapor stream and liquid stream.

The vapor stream is mainly EL, PL, and M and is recycled back.

Some EL will leave in the vapor stream which can be recovered via an absorber.

The Synthesis Step: From Ill-posed Problems to PFD – p.29/32

Increasing the purge stream results in loss of reactants but reduces the amount of methane going into the reactor (M has to less than 10% to prevent coking).

The liquid recovery system is designed as a series of distillation columns.

The Synthesis Step: From Ill-posed Problems to PFD – p.30/32

Increasing the purge stream results in loss of reactants but reduces the amount of methane going into the reactor (M has to less than 10% to prevent coking).

The liquid recovery system is designed as a series of distillation columns.

The Synthesis Step: From Ill-posed Problems to PFD – p.30/32

Increasing the purge stream results in loss of reactants but reduces the amount of methane going into the reactor (M has to less than 10% to prevent coking).

The liquid recovery system is designed as a series of distillation columns.

The Synthesis Step: From Ill-posed Problems to PFD – p.30/32

Final Design

Distillation is carried out in order of decreasing volatility of species.

The Synthesis Step: From Ill-posed Problems to PFD – p.31/32

Final Design

Distillation is carried out in order of decreasing volatility of species.

The Synthesis Step: From Ill-posed Problems to PFD – p.31/32

Final Design

Distillation is carried out in order of decreasing volatility of species.

The Synthesis Step: From Ill-posed Problems to PFD – p.31/32

Every flowsheet has hot streams that need cooling and cold streams that require heating.

Option 1: Use separate utility streams for each heating/cooling process stream. Advantage: Easy to set-up and operate. Disadvantage: Expensive to build and maintain.

Option 2: Exchange energy between process streams. Advantage: Saves cost. Disadvantage: More complex process, difficult to control.

Which option to use depends on process economics.

Every flowsheet has hot streams that need cooling and cold streams that require heating.

Option 1: Use separate utility streams for each heating/cooling process stream. Advantage: Easy to set-up and operate. Disadvantage: Expensive to build and maintain.

Option 2: Exchange energy between process streams. Advantage: Saves cost. Disadvantage: More complex process, difficult to control.

Which option to use depends on process economics.

Every flowsheet has hot streams that need cooling and cold streams that require heating.

Option 1: Use separate utility streams for each heating/cooling process stream. Advantage: Easy to set-up and operate. Disadvantage: Expensive to build and maintain.

Option 2: Exchange energy between process streams. Advantage: Saves cost. Disadvantage: More complex process, difficult to control.

Which option to use depends on process economics.

Every flowsheet has hot streams that need cooling and cold streams that require heating.

Option 1: Use separate utility streams for each heating/cooling process stream. Advantage: Easy to set-up and operate. Disadvantage: Expensive to build and maintain.

Option 2: Exchange energy between process streams. Advantage: Saves cost. Disadvantage: More complex process, difficult to control.

Which option to use depends on process economics.

Every flowsheet has hot streams that need cooling and cold streams that require heating.

Option 1: Use separate utility streams for each heating/cooling process stream. Advantage: Easy to set-up and operate. Disadvantage: Expensive to build and maintain.

Option 2: Exchange energy between process streams. Advantage: Saves cost. Disadvantage: More complex process, difficult to control.

Which option to use depends on process economics.

Every flowsheet has hot streams that need cooling and cold streams that require heating.

Option 1: Use separate utility streams for each heating/cooling process stream. Advantage: Easy to set-up and operate. Disadvantage: Expensive to build and maintain.

Option 2: Exchange energy between process streams. Advantage: Saves cost. Disadvantage: More complex process, difficult to control.

Which option to use depends on process economics.

✓Gather information about the process chemistry ✓Generate flow diagram based on Douglas Hierarchy Solve mass and energy balances Estimate equipment size based on flow rates from previous step Estimate equipment cost based on size from previous step Optimize process